# Progressive Token Length Scaling in Transformer Encoders for Efficient Universal Segmentation

### v1

@@ 3.1 PRO-SCALE: Proposed Transformer Encoder
The bottleneck towards an efficient encoder are excessive large-scale features, where most of which are not informative but contain local details for different objects [18]. In order to better handle their usage in the encoder, we can leverage the composition of multi-scale features from the backbone: (1) The limited quantity of small-scale features (e.g. ${\bm{s}}_{3}$, ${\bm{s}}_{4}$) captures abundant semantics. (2) The larger quantity of large-scale features (e.g. ${\bm{s}}_{1}$, ${\bm{s}}_{2}$) captures crucial local features important for segmenting varying scale of objects. For example, assuming a SWIN-T backbone based M2F, ${\bm{s}}_{2},{\bm{s}}_{3}$, ${\bm{s}}_{4}\}$ result in {76.19%, 19.04%, 4.76%} token contributions, respectively. Therefore, we propose to prioritize enrichment of tokens from small scale features earlier than the large scale features.
@@
We build upon deformable attention as it exhibits linear complexity with the number of feature queries (in our case, represented by ${\bm{P}}$). We create three splits of ${\bm{P}}$ (following the three scales):
${\bm{P}}_{1}$ $=\mathcal{C}({\bm{s}}_{4})\in\mathbb{R}^{K_{1}\times C},$ (1)
${\bm{P}}_{2}$ $=\mathcal{C}({\bm{s}}^{\prime},{\bm{s}}_{3})\in\mathbb{R}^{K_{2}\times C}$ (2)
${\bm{P}}_{3}$ $=\mathcal{C}({\bm{s}}^{\prime\prime},{\bm{s}}_{2})\in\mathbb{R}^{K_{3}\times C}$ (3)
Here, $\mathcal{C}(\cdot)$ denotes the concatenation operation along the token size dimension, $K_{3}=\sum(\nicefrac{{HW}}{{32^{2}}}+\nicefrac{{HW}}{{16^{2}}}+\nicefrac{{HW}}{{8^{2}}}),K_{2}=\sum(\nicefrac{{HW}}{{32^{2}}}+\nicefrac{{HW}}{{16^{2}}})$, and $K_{1}=\nicefrac{{HW}}{{32^{2}}}$. These splits are then fed to three stages of the deformable attention transformer layers sequentially. The output of these stages are ${\bm{s}}^{\prime},{\bm{s}}^{\prime\prime}$, and ${\bm{s}}^{\prime\prime\prime}$, respectively. ${\bm{s}}^{\prime\prime\prime}$ is subsequently fed to the decoder. Each stage is repeated $p_{1},p_{2}$ and $p_{3}$ times before propagating the tokens to the next stage. This results in updation of ${\bm{s}}_{4}$ for $(p_{1}+p_{2}+p_{3})$, ${\bm{s}}_{3}$ for $(p_{2}+p_{3})$, and ${\bm{s}}_{2}$ for $p_{1}$ times, gradually expanding the token length of inputs to intermediate layers in PRO-SCALE. As we show in Sec. 4, this strategy reduces computing load by significant margins while maintaining performance.
Figure 4: Token Re-Calibration (TRC) module. This unit first interpolates ${\bm{s}}_{2}$ to the size of the target smaller-scale feature ${\bm{s}}_{i}$. Then, the spatially flattened and resized ${\bm{s}}_{2}$ is passed through a channel reduction layer $\phi$ (followed by sigmoid function) to produce an attention map. This map is imposed on the spatially flattened ${\bm{s}}_{i}$ to create $\widehat{{\bm{s}}}_{i}$.
3.1.1 Token Re-Calibration (TRC) Module
Intuition.
Skipping the propagation of ${\bm{s}}_{2}$ in the initial layers may lead to errors in predicted panoptic maps due to missing information from this feature scale. To avoid such map errors caused by imperfect token representations, we propose to calibrate the ${\bm{s}}_{3}$ and ${\bm{s}}_{4}$ using ${\bm{s}}_{2}$ in our TRC module. This module aims to enhance small-scale features ${\bm{s}}_{i}\in\{{\bm{s}}_{3},{\bm{s}}_{4}\}$ using large-scale feature ${\bm{s}}_{2}$ without increasing the computations in PRO-SCALE transformer layers.
TRC structure.
To utilize strengths of ${\bm{s}}_{2}$ in smaller scales, we employ contrastive attention. In particular, we propose to enrich the tokens of ${\bm{s}}_{i}$ by projecting ${\bm{s}}_{2}$ on ${\bm{s}}_{i}$ via an attention map. This attention map is obtained by using a channel reduction layer $\phi$ followed by a sigmoid function. As illustrated in Fig. 4, the final tokens $\widehat{{\bm{s}}}_{i}$ are obtained as ($\otimes$ represents elementwise multiplication):
$\widehat{{\bm{s}}}_{i}={\bm{s}}_{i}\otimes{\texttt{sigmoid}(\epsilon\phi({\bm{s}}_{2}))}$ (4)
With above, Eq. 2, 3 translate to ${\bm{P}}_{1}=\mathcal{C}(\widehat{{\bm{s}}}_{4})$ and ${\bm{P}}_{2}=\mathcal{C}({\bm{s}}^{\prime},\widehat{{\bm{s}}}_{3})$. We choose $\phi(\cdot)$ to be a linear layer and set temperature $\epsilon=0.1$, keeping computations negligible.
@@ 3.1.2 Light-Pixel Embedding (LPE) Module
Strong performance of M2F depends on multi-scale features computed from the backbone [5]. Tokens $\{{\bm{s}}_{2},{\bm{s}}_{3},{\bm{s}}_{4}\}$ are fed to the encoder layers to compute ${\bm{s}}^{\prime\prime\prime}$ in order to produce per-segment embeddings in the transformer decoder. ${\bm{s}}_{1}$, on the other hand, serves the purpose of creating the per-pixel embedding map $\mathcal{E}_{emb}$. However, it results in high computational load due to use of convolutional layer in original M2F. Hence rather than dropping ${\bm{s}}_{1}$, we assess the existing learnable module with a simpler module in our design.
@@
We propose to use a simple maxpooling layer followed by normalization and non-linearity to compute $\mathcal{E}_{emb}$. The goal of our LPE module is to mitigate this overhead while keeping the advantages of ${\bm{s}}_{1}$ in producing $\mathcal{E}_{emb}$. We observe it reduces almost $\sim$45GFLOPs in the encoder architecture, but doesn’t significantly harm the overall model’s segmentation performance. In our implementation, we use a pooling kernel size of 3.
@@ 4 Experiments
In this section, we evaluate PRO-SCALE based M2F architecture on two benchmarks on all segmentation tasks. First, we observe that PRO-SCALE is extremely effective in reducing computational load while providing best trade-offs compared to baselines (Tab. 1, 2). Second, we analyze each additional component of PRO-SCALE (TRC and LPE units) in Tab. 4, 5, and 8, and Fig. 6. Third, we analyze the impact of different backbones and their pre-training strategies in Tab. 6 and Fig. 8. Finally, we visualize some segmentation examples in Fig. 9. In our Supplementary material, we provide our training, dataset (for COCO [21] and Cityscapes [32]), and baseline details. We also provide additional results on larger backbones, analysis with object detection as well as open-vocab segmentation which further underscore the strengths of our encoder. For brevity, we will denote our overall framework as PRO-SCALE.
@@
We evaluate PRO-SCALE and baselines in the universal segmentation setting. This means, following [5], PRO-SCALE’s evaluation is performed using a model trained exclusively with panoptic segmentation annotations. For panoptic segmentation, the conventional PQ (Panoptic Quality [3]) metric is used. Following [5], we report APp (Average-Precision [21]) metric for instance segmentation. This is computed on the ‘thing’ categories from the instance segmentation annotations. For semantic segmentation, we report mIoUp (mean Intersection-over-Union [33]) by merging instance masks from the same category. Here, subscript $p$ denotes evaluation after training with panoptic segmentation annotations. For computing the GFLOPs, we use image scale of (800, 1333) for COCO dataset, and (1024, 2048) for Cityscapes dataset. All models are trained and evaluated on the train and validation split, respectively.
@@
We focus on standard lightweight backbones Res50 [16], SWIN-Tiny [17], and MViT2-Tiny [34], pre-trained on ImageNet-1K [35]. We use PRO-SCALE as the transformer encoder. As per Sec. 3, we use different integer values of ($p_{1},p_{2}$, $p_{3}$) to instantiate different depths of the stages in PRO-SCALE. We directly adopt the M2F’s decoder that consists of masked attention [5] with 9 layers in total and 100 learnable queries. Similar to M2F’s round robin design, PRO-SCALE feeds $\{{\bm{s}}_{2},{\bm{s}}_{3},{\bm{s}}_{4}\}$ into successive transformer decoder layers.
Table 1: COCO evaluation. PRO-SCALE is extremely competitive against the baselines on COCO with at least 51.99% GFLOPs reduction compared to M2F with no performance drop. Non-M2F-style architectures are colored in gray.
@@
Model PQ mIOUp APp Total Tx. Enc.
@@
M2F [5] 51.73 61.94 41.72 229.10 135.00
MF [4] 46.50 57.80 33.00 181.00 (-20.99) – (–)
PEM [8] 46.38 55.95 34.25 110.90 (-51.59) – (–)
YOSO [13] 48.40 58.74 36.87 114.50 (-50.02) – (–)
RAP-SAM [30] 46.90 123.00 (-46.31) – (–)
ReMax [15] 53.50 – (–) – (–)
@@
PRO-SCALE (2, 2, 2) 51.21 61.53 40.66 148.8 (-35.05) 48.85 (-63.48)
PRO-SCALE (3, 3, 3) 51.45 61.58 41.45 166.1 (-27.50) 67.45 (-50.03)
Table 2: Cityscapes evaluation. PRO-SCALE shows strong efficiency trade-off compared to the baselines on Cityscapes, e.g. 51.96% (SWIN-T) and 50.17% (Res50) GFLOPs reduction and little-to-no accuracy drop. Non-M2F-style models are colored in gray. $\dagger$ trained for 200K iterations.
@@
Model PQ IoUp APp Total Tx. Enc.
@@
M2F [5] 64.00 80.77 39.26 537.8 281.00
Lite-M2F [18] 62.29 79.43 36.57 428.7 (-20.29) 172.00 (-38.79)
RT-M2F [19] 59.73 77.89 31.35 361.1 (-32.86) 130.00 (-53.74)
PRO-SCALE (1, 1, 1) 60.58 78.29 32.97 311.9 (-42.00) 055.14 (-80.38)
PRO-SCALE (2, 2, 2) 62.37 78.62 35.97 352.0 (-34.55) 095.24 (-66.11)
PRO-SCALE (3, 3 ,3) 63.06 77.94 37.81 392.1 (-27.09) 135.00 (-51.96)

### v2

@@ 3.1 PRO-SCALE: Proposed Transformer Encoder
The bottleneck towards an efficient encoder are excessive large-scale features, where most of which are not informative but contain local details for different objects (Li et al., 2023a). In order to better handle their usage in the encoder, we can leverage the composition of multi-scale features from the backbone: (1) The limited quantity of small-scale features (e.g. ${\mathbf{s}}_{3}$, ${\mathbf{s}}_{4}$) captures abundant semantics. (2) The larger quantity of large-scale features (e.g. ${\mathbf{s}}_{1}$, ${\mathbf{s}}_{2}$) captures crucial local features important for segmenting varying scale of objects. For example, assuming a SWIN-T backbone based M2F, ${\mathbf{s}}_{2},{\mathbf{s}}_{3}$, ${\mathbf{s}}_{4}\}$ result in {76.19%, 19.04%, 4.76%} token contributions, respectively (also see our analysis in Fig. 9 in Appendix). Therefore, we propose to prioritize enrichment of tokens from small scale features earlier than the large scale features.
@@
We build upon deformable attention as it exhibits linear complexity with the number of feature queries (in our case, represented by ${\mathbf{P}}$). We create three splits of ${\mathbf{P}}$ (following the three scales):
${\mathbf{P}}_{1}$ $=\mathcal{C}({\mathbf{s}}_{4})\in\mathbb{R}^{K_{1}\times C},$ (1)
${\mathbf{P}}_{2}$ $=\mathcal{C}({\mathbf{s}}^{\prime},{\mathbf{s}}_{3})\in\mathbb{R}^{K_{2}\times C}$ (2)
${\mathbf{P}}_{3}$ $=\mathcal{C}({\mathbf{s}}^{\prime\prime},{\mathbf{s}}_{2})\in\mathbb{R}^{K_{3}\times C}$ (3)
Here, $\mathcal{C}(\cdot)$ denotes the concatenation operation along the token size dimension, $K_{3}=\sum(\nicefrac{{HW}}{{32^{2}}}+\nicefrac{{HW}}{{16^{2}}}+\nicefrac{{HW}}{{8^{2}}}),K_{2}=\sum(\nicefrac{{HW}}{{32^{2}}}+\nicefrac{{HW}}{{16^{2}}})$, and $K_{1}=\nicefrac{{HW}}{{32^{2}}}$. These splits are then fed to three stages of the deformable attention transformer layers sequentially. The output of these stages are ${\mathbf{s}}^{\prime},{\mathbf{s}}^{\prime\prime}$, and ${\mathbf{s}}^{\prime\prime\prime}$, respectively. ${\mathbf{s}}^{\prime\prime\prime}$ is subsequently fed to the decoder. Each stage is repeated $p_{1},p_{2}$ and $p_{3}$ times before propagating the tokens to the next stage. This results in updation of ${\mathbf{s}}_{4}$ for $(p_{1}+p_{2}+p_{3})$, ${\mathbf{s}}_{3}$ for $(p_{2}+p_{3})$, and ${\mathbf{s}}_{2}$ for $p_{1}$ times, gradually expanding the token length of inputs to intermediate layers in PRO-SCALE. As we show in Sec. 4, this strategy reduces computing load by significant margins while maintaining performance. Note that, since we skip using tokens from large scale features in the initial layers, we also use a token recalibration within PRO-SCALE that enriches small-scale features with high-scale features to gain slightly better segmentation without significant computational overhead. Please see Sec. C in Appendix for more details and Fig. 7 for a complete visualization.
@@ 3.1.1 Light-Pixel Embedding (LPE) Module
Strong performance of M2F depends on multi-scale features computed from the backbone (Cheng et al., 2022). Tokens $\{{\mathbf{s}}_{2},{\mathbf{s}}_{3},{\mathbf{s}}_{4}\}$ are fed to the encoder layers to compute ${\mathbf{s}}^{\prime\prime\prime}$ in order to produce per-segment embeddings in the transformer decoder. ${\mathbf{s}}_{1}$, on the other hand, serves the purpose of creating the per-pixel embedding map $\mathcal{E}_{emb}$ to enhance local details in the feature maps. However, due to the large size of ${\mathbf{s}}_{1}$, it results in high computational load from the use of convolutional layers in original M2F. Hence rather than dropping ${\mathbf{s}}_{1}$, we weaken this inductive bias and assess the existing learnable module with a simpler module in our design.
@@
We propose to use a simple maxpooling layer followed by normalization and non-linearity to compute $\mathcal{E}_{emb}$. The goal of our LPE module is to mitigate this overhead while keeping the advantages of ${\mathbf{s}}_{1}$ in producing $\mathcal{E}_{emb}$. We observe it reduces almost $\sim$45GFLOPs in the encoder architecture, but doesn’t significantly harm the overall model’s segmentation performance. In our implementation, we use a pooling kernel size of 3 ( and stride is set to 1).
@@ 4 Experiments
In this section, we evaluate PRO-SCALE based M2F architecture on two benchmarks on all segmentation tasks. First, we observe that PRO-SCALE is extremely effective in reducing computational load while providing best trade-offs compared to baselines (Tab. 1, 2, 11). Second, we provide an extensive ablation analysis of PRO-SCALE in Tab. 3 - 7, and Fig. 5, 5. Third, we test the performane of PRO-SCALE on prevalent state-of-the-art frameworks for tasks like object detection, joint-task predictions, and open-vocabulary universal segmentation in Tab. 11, 11 and 11. Finally, we visualize some segmentation examples in Fig. 6. In the Appendix, we provide our training, and dataset (for COCO (Lin et al., 2014) and Cityscapes (Cordts et al., 2016)) details. For brevity, we will denote our overall framework as PRO-SCALE.
@@
We evaluate PRO-SCALE and baselines in the universal segmentation setting. This means, following (Cheng et al., 2022), PRO-SCALE’s evaluation is performed using a model trained exclusively with panoptic segmentation annotations. For panoptic segmentation, the conventional PQ (Panoptic Quality (Kirillov et al., 2019)) metric is used. Following (Cheng et al., 2022), we report APp (Average-Precision (Lin et al., 2014)) metric for instance segmentation. This is computed on the ‘thing’ categories from the instance segmentation annotations. For semantic segmentation, we report mIoUp (mean Intersection-over-Union (Everingham et al., 2015)) by merging instance masks from the same category. Here, subscript $p$ denotes evaluation after training with panoptic segmentation annotations. For computing the GFLOPs, we use image scale of (800, 1333) for COCO dataset, and (1024, 2048) for Cityscapes dataset. All models are trained and evaluated on the train and validation split, respectively.
Baselines.
We compare PRO-SCALE with the original M2F, along with some recently proposed efficient transformer encoders for detection (Li et al., 2023a; Lv et al., 2023; Cavagnero et al., 2024). Specifically, we replace the transformer encoder of M2F with encoders proposed in Lite-DETR (Li et al., 2023a) and RT-DETR (Lv et al., 2023). We call these Lite-M2F and RT-M2F, respectively. Note that, we opted for the “Lite-DETR H3L1-(6+1)$\times$1” setup without its key-aware deformable attention (Li et al., 2023a). However, we modified this setup to (5+1) when implementing Lite-M2F. For completeness, we also compare with recent non-M2F-style panoptic segmentation models YOSO (Hu et al., 2023), RAP-SAM (Xu et al., 2024) and ReMaX (Sun et al., 2023).
@@
We focus on standard lightweight backbones Res50 (He et al., 2016), SWIN-Tiny (Liu et al., 2021), and MViT2-Tiny (Li et al., 2022b), pre-trained on ImageNet-1K (Deng et al., 2009). We use PRO-SCALE as the transformer encoder. As per Sec. 3, we use different integer values of ($p_{1},p_{2}$, $p_{3}$) to instantiate different depths of the stages in PRO-SCALE. We directly adopt the M2F’s decoder that consists of masked attention (Cheng et al., 2022) with 9 layers in total and 100 learnable queries. Similar to M2F’s round robin design, PRO-SCALE feeds $\{{\mathbf{s}}_{2},{\mathbf{s}}_{3},{\mathbf{s}}_{4}\}$ into successive transformer decoder layers.
Table 1: COCO evaluation. PRO-SCALE (with configuration ($p_{1}$, $p_{2}$, $p_{3}$)) is extremely competitive against the baselines on COCO with at least 51.99% GFLOPs reduction compared to M2F with no performance drop. Non-M2F-style architectures are colored in gray.
@@
Model PQ mIOUp APp Total Encoder
@@
M2F (Cheng et al., 2022) 51.73 61.94 41.72 229.10 135.00
MF (Cheng et al., 2021) 46.50 57.80 33.00 181.00 (-20.99) – (–)
PEM (Cavagnero et al., 2024) 46.38 55.95 34.25 110.90 (-51.59) – (–)
YOSO (Hu et al., 2023) 48.40 58.74 36.87 114.50 (-50.02) – (–)
RAP-SAM (Xu et al., 2024) 46.90 123.00 (-46.31) – (–)
ReMaX (Sun et al., 2023) 53.50 169.00 (-26.23) – (–)
@@
PRO-SCALE (2, 2, 2) 51.21 61.53 40.66 148.80 (-35.05) 48.85 (-63.48)
PRO-SCALE (3, 3, 3) 51.45 61.58 41.45 166.10 (-27.50) 67.45 (-50.03)
Table 2: Cityscapes evaluation. PRO-SCALE (with configuration ($p_{1}$, $p_{2}$, $p_{3}$)) shows strong efficiency trade-off compared to the baselines on Cityscapes, e.g. 51.96% (SWIN-T) and 50.17% (Res50) GFLOPs reduction and little-to-no accuracy drop. Non-M2F-style models are colored in gray. $\dagger$ trained for 200K iterations.
@@
Model PQ IoUp APp Total Encoder
@@
M2F (Cheng et al., 2022) 64.00 80.77 39.26 537.8 0 281.00
Lite-M2F (Li et al., 2023a) 62.29 79.43 36.57 428.70 (-20.29) 172.00 (-38.79)
RT-M2F (Lv et al., 2023) 59.73 77.89 31.35 361.10 (-32.86) 130.00 (-53.74)
PRO-SCALE (1, 1, 1) 60.58 78.29 32.97 311.90 (-42.00) 055.14 (-80.38)
PRO-SCALE (2, 2, 2) 62.37 78.62 35.97 352.00 (-34.55) 095.24 (-66.11)
PRO-SCALE (3, 3 ,3) 63.06 77.94 37.81 392.10 (-27.09) 135.00 (-51.96)